### v1

# SurfDist: Interpretable Three-Dimensional Instance Segmentation Using Curved Surface Patches

Abstract
We present SurfDist, a convolutional neural network architecture for three-dimensional volumetric instance segmentation. SurfDist enables prediction of instances represented as closed surfaces composed of smooth parametric surface patches, specifically bicubic Bézier triangles. SurfDist is a modification of the popular model architecture StarDist-3D which breaks StarDist-3D’s coupling of instance parameterization dimension and instance voxel resolution, and it produces predictions which may be upsampled to arbitrarily high resolutions without introduction of voxelization artifacts.
For datasets with blob-shaped instances, common in biomedical imaging, SurfDist can outperform StarDist-3D with more compact instance parameterizations. We detail SurfDist’s technical implementation and show one synthetic and one real-world dataset for which it outperforms StarDist-3D. These results demonstrate that interpretable instance surface models can be learned effectively alongside instance membership.
1 Background
In image segmentation, machine learning approaches have achieved major advances over classical computer vision techniques. However, these approaches pose mechanistic interpretability challenges that can make it difficult both to reason about a model’s ability to generalize beyond its training dataset and to value synthetic data generated by a trained model. Additionally, instance segmentation of 3D volumetric data remains undeveloped compared to instance segmentation of 2D images, because of additional computational expense associated with 3D machine learning models as well as a lack of publicly available high-quality 3D benchmark datasets.
The StarDist family of models has become popular for biomedical image segmentation tasks where instances are blob-like objects. These models are constructed by augmenting "backbone" instance segmentation networks, such as U-Net ronneberger2015u , with additional output layers which enforce instance embeddings into an explicitly parameterized geometric space. This construction both aids interpretability and improves segmentation accuracy relative to backbone networks alone.
In the original StarDist model schmidt2018cell , instances within 2D images are modeled as star-convex polygons (Figure 1(a)). This parameterization often works well in practice, but because it predicts piecewise linear approximations of instance boundaries, its ability to model local curvature degrades as instance sizes (in pixels) increase. The SplineDist model mandal2021splinedist solves this problem by modifying StarDist to model instances as closed spline curves (Figure 1(b)) rather than as polygons. This approach decouples model complexity from instance spatial resolution, and empirically, it reduces, relative to StarDist, the number of network parameters required to achieve a given level of segmentation accuracy.
StarDist-3D weigert2020star extends the 2D StarDist approach to the 3D setting by modeling instances as star-convex polyhedra. Like the 2D StarDist model, it works well for many current datasets but couples local curvature modeling to instance spatial extent. Toward breaking this coupling in the 3D setting, we present SurfDist, a modification of StarDist-3D which models instances as contiguous, closed meshes of bicubic triangular surface patches. We find that SurfDist achieves segmentation accuracy comparable to or better than that of StarDist-3D on real biological data while fitting simpler instance representations.
2 Technical details
2.1 Mesh parameterization
SurfDist models instances as meshes of curved triangular surface patches. Each patch is formulated as a bicubic Bézier triangle parameterized by 10 control points (Figure 2). Of these 10 control points, 3 are at the triangle’s vertices (and are guaranteed to be on the resulting surface), 6 are associated with (but generally not on) the triangle’s edges, and 1 is associated with (but generally not on) the triangle’s interior. Points on the patch surface are given by the bicubic interpolation
across a barycentric triangular input domain where is the set of control points as indexed in Figure 2 (alma991085856964006532 ).
Since SurfDist meshes comprise multiple trianglular patches, and since adjacent patches share vertices and edges, all control points except those associated with triangle interiors are shared by multiple triangles in a given mesh, and the total number of control points for a given mesh is given by where is the number of vertices, is the number of edges, and is the number of triangles in the mesh. As in StarDist-3D, is the only value of these three which is a specifiable hyperparameter. and are determined by the triangular mesh given by taking the convex hull of a mapping of each vertex in to a position on the unit sphere using either the Fibonacci lattice gonzalez2010measurement or an equidistant spacing scheme.
Once a triangular mesh has been calculated, the control points for each triangle are assumed to lie along the rays drawn from the unit sphere’s center toward the triangle’s vertices (for vertex control points), toward its center (for interior control points), or toward points that divide its edges into thirds (for edge control points). Instance predictions then simply comprise radial distance estimates for each control point, so the number of free parameters for a given instance is equal to the number of control points in its mesh (Figure 3(b)). This construction results in a topologically closed mesh which is smooth across triangular surfaces and continuous across the edges of constituent triangles. Importantly, mesh surfaces and edges may be iteratively subdivided (kato2002curved ). This property enables direct use of StarDist-3D’s inference implementation (by using subdivision to approximate curved meshes with arbitrarily high-detailed polyhedra), and it allows for sampling of surface points at arbitrarily fine resolutions during both model training and inference.
2.2 Loss
Loss within StarDist family models is formulated as a sum of object () and distance () terms. The object term is a probability, and the distance term has absolute magnitude. These terms are defined at the voxel level: for a given voxel, the object term tracks foreground identity confidence, and the distance term tracks predicted instance fit. Loss for a given input image is calculated by taking the mean absolute error (MAE) of loss across all of its voxels. Formally, per-voxel loss is defined as
where and are actual and predicted distances to the nearest exterior (either background or interior of another instance) voxel normalized by the maximum such distance for the voxel’s instance, is standard binary cross-entropy loss, is a specifiable regularization weight, and are vectors of predicted and actual radial distances to instance exterior along radial directions, and
where is another specifiable regularization weight.
These loss equations are identical to the ones used in StarDist-3D (weigert2020star ), but there are two important differences between StarDist-3D’s and SurfDist’s loss formulations related to the set of radial directions . First, in StarDist-3D, is determined programatically at model instantiation time using a spherical Fibonacci lattice (gonzalez2010measurement ) and observed instance anisotropy across the training dataset, and it is then held constant across all voxels for all training and inference steps. In SurfDist, is determined on a per-voxel and per-prediction basis as
where is a set of surface points sampled from the smooth triangular faces of the mesh predicted by the model for a given voxel and is the coordinate vector of that voxel.
Second, in StarDist-3D, the number of rays used to calculate the term is defined by the number of parameters in the output layer of the model. In SurfDist, is determined by the surface point sampling paradigm (which is a hyperparameter), and increasing or decreasing the value of does not change the parameter count of the trained model (although it will affect memory requirements for model training steps). This means that even low parameter-count SurfDist models may be trained to predict entire instance surfaces (rather than just the subset of surface points given by radial boundary intersections which would be predicted by a StarDist-3D model). In our experiments, we draw equidistant barycentric samples from the smooth triangular faces of all predicted instances by evaluating the barycentric points resulting from two successive subdivisions of a flat triangle. We find that this sampling paradigm provides a reasonable balance between morphological detail and resource requirements during model training.
3 Experimental results
We present results for one synthetic toy dataset, Sphere, and one publicly available volumetric microscopy dataset, Worm long20093d .
3.1 Reconstruction of Synthetic Spheres
Sphere was constructed to trivially validate SurfDist’s modeling approach by illustrating its ability to model objects with smooth surfaces at higher fidelity and with fewer parameters than StarDist-3D. Each of its images is a single centered spherical mask voxel set (generated by scikit-image ’s morphology.ball function) with spherical radial lengths varying across images. For each image, we compared reconstruction accuracy between SurfDist and StarDist-3D instance parameterizations across a range of instance ray counts. For StarDist-3D reconstructions, radial distances from image centers were measured directly. For SurfDist reconstructions, vertex control point distances were set using spherical radial lengths, and face and edge control point distances (which were assumed to be the same across all face and across all edge control points for a given image) were optimized using 2020SciPy-NMeth ’s optimize.least_squares. Our results (Figure 4) show that as instance size increases, SurfDist’s parameterization is much better at reconstructing spherical input than StarDist-3D’s. Generally, this is true even when SurfDist is allowed many fewer parameters than StarDist-3D.
3.2 Segmentation of Real Volumetric Microscopy Data
Worm was chosen as a real-data benchmark because it was one of the datasets used as a benchmark in weigert2020star . We trained five StarDist-3D and ten SurfDist models with an 18/3 train/val split. To match the recommended default parameterization for StarDist-3D and the parameterization for which results on Worm were reported in weigert2020star , we used a 96-ray StarDist-3D model for all five StarDist-3D training runs. For SurfDist, to demonstrate its relative compactness, we trained five models each for two parameterizations: one using 6 rays (38 total free parameters) per instance and one using 12 rays (92 total free parameters) per instance. Comparison of validation segmentation accuracy of trained models (Table 1) shows that 6-ray SurfDist models performed comparably to 96-ray StarDist-3D models on average, and our best 6-ray SurfDist model outperformed all 96-ray StarDist-3D models. 12-ray SurfDist models outperformed 96-ray StarDist-3D models both on average and in the best case. Notably, instance resolution in this dataset is relatively low, a feature which should minimize SurfDist’s advantage over StarDist-3D. Example renderings of meshes predicted for this dataset by one of our trained SurfDist models are shown in Figure 6.
Average Metrics Across All Trained Models by Model Type | |||||
Model | Precision | Recall | Accuracy | F1 | Panoptic Quality |
SurfDist 6 | 0.6141 | 0.6147 | 0.5353 | 0.6144 | 0.4068 |
SurfDist 12 | 0.6221 | 0.6248 | 0.5430 | 0.6234 | 0.4185 |
StarDist-3D 96 | 0.6141 | 0.6069 | 0.5285 | 0.6105 | 0.4047 |
Average Metrics for Best Model by Model Type | |||||
Model | Precision | Recall | Accuracy | F1 | Panoptic Quality |
SurfDist 6 | 0.6320 | 0.6343 | 0.5518 | 0.6331 | 0.4303 |
SurfDist 12 | 0.6245 | 0.6327 | 0.5466 | 0.6285 | 0.4253 |
StarDist-3D 96 | 0.6254 | 0.6117 | 0.5365 | 0.6185 | 0.4147 |
4 Discussion
SurfDist learns representations of three-dimensional instances as closed meshes of parameterized curved surface patches. Experimental results demonstrate that it can achieve accuracy superior to that of StarDist-3D with a smaller number of parameters for blob-shaped objects which are common in biomedical imaging data. SurfDist’s experimental success shows that the StarDist architectural approach is effective for training machine learning models to estimate parametric surfaces defining 3D blob-like data, and it motivates further development of computational geometry approaches for three-dimensional instance segmentation.
Despite its state-of-the-art performance on segmentation tasks, SurfDist has two significant limitations which should inform the direction of future work. The first of these is that, like StarDist-3D, SurfDist is ill-suited for datasets where instances are not star-convex, such as branching dendritic cells. Second, additional work is needed to reduce the computational time required for inference using trained SurfDist (or StarDist-3D) models. The bottleneck in inference is the non-maximum suppression step, in which an instance is predicted for every voxel of the input volume and pairwise intersections of predicted instances are computed to de-duplicate predictions. Computation of intersections of objects bounded by polyhedral meshes is a common task in computer graphics and already highly optimized, so a fruitful avenue for reducing inference costs may be development of alternative schemes to non-maximum suppression which decouple instance predictions from input voxel grids.
5 Data and code availability
Code for the SurfDist model and for reproduction of the experiments detailed here is available at https://github.com/focolab/SurfDist.